If I'm performing GPC analysis of Sample A in Sample B at a given temperature and I'm given elution volume and concentration values. I then need to get those values into weights. When I do this will I use the MW of Sample A or Sample B to do this? Equation for reference below:

If my thinking of any of the following is also wrong, please inform me, thanks!

(Concentration)*(Elution Volume)*(Molecular Weight) = Weight

(Mol/L)*(L)*(g/mol) = g
Answer #1

Let us have a mL of x M sample A mixed with b mL of y M sample B.

The total volume of the sample = (a + b) mL which is diluted by the solvent to get the elution volume, say V mL.

The concentrations of samples A and B in the eluted volume can be easily found out as the final concentration = (initial concentration)*(initial volume)/(final volume)

Therefore,

[A] = (x M)*(a mL)/(V mL)

= ax/V mM (mM = millimolar)

[B] = (y M)*(b Ml)/(V mL)

= by/V mM

Number of moles of A = (final concentration of A)*(final volume)

= (ax/V mM)*(V mL)

= ax mol

Number of moles of B = (final concentration of B)*(final volume)

= (by/V mM)*(V mL)

= by mol

Let MA and MB g/mol be the molar masses of samples A and B.

Mass or weight of A in the GPC mixture = (ax mol)*(MA g/mol)

= axMA g.

Mass or weight of B in the GPC mixture = (by mol)*(MB g/mol)

= byMB g (ans).